# SkyAI Sim: An Open-Source Simulation of UAV Aerial Imaging from Satellite Data

### v1

@@ 3.3.1 Input Formats for the Argument Parser
    The default is "SkyAI". It can be switched to "VBN" to access and process the Supplementary data.
@@
    vmargin: The vertical margin is an optional but crucial parameter. Google permits publishing images with watermarks, but these can bias data. To avoid this, we capture images with an extra vertical margin—20% total, split as 10% at the top and 10% at the bottom. This margin is later cropped out during processing by the "preprocess_image" method in the Google Maps module.
@@ 3.4 Post Processing
To cleanup the data as an optional step, a threshold entropy is defined. The threshold is selected according to the distribution of entropies in all images, along with experimental observations provided in Figure 6. Specifically in the Agricenter dataset, the images below an entropy of 2.1 lack significant details. To confirm this value’s reliability and ensure it doesn’t filter out substantial valuable data, a barplot and histogram of entropies for all available roadmap images are shown in Figures 10 and 11. This ensures most data is above the threshold, making it safe to discard images below $2.1$. SkyAI performs data cleaning based on the map type. If the map type is "satellite" or anything other than "roadmap", it first checks for "roadmap" data in the data directory. If available, images are filtered according to "roadmap" images with an entropy above the default threshold (2.1). If no roadmap data exists, images are filtered based on their original map type.
@@ 4 Discussion
SkyAI Sim opens the door to numerous applications, but it does have its limitations. Some of these are inherited from the API it relies on. A notable limitation is the inability to determine the exact date of the imagery captured through the Google Maps Static API. Google does not provide this information via the API, and the dates can vary widely depending on the region and zoom level. This is due to Google Maps using a mix of satellite, aerial, and Street View imagery from different times. Since the Google Maps Static API does not provide the date of image capture, this limitation, while not critical for most of SkyAI Sim’s potential applications, is still an important area for future improvement. One possible solution is to integrate additional data sources that offer time-stamped satellite imagery. For instance, APIs from services like Google Earth Engine [7] could be leveraged to retrieve historical satellite images with precise dates. This would enhance the accuracy and utility of static map images in applications where temporal information is crucial.
Another limitation is the restriction on the number of images that can be freely published. If SkyAI Sim is used to generate and publish a specific dataset, no more than 5,000 images can be released without obtaining permission from Google. Additionally, free access to data retrieval is not unlimited; monthly restrictions apply to the amount of data that can be retrieved for free. However, these limitations are manageable, as obtaining permission or making additional payments can easily resolve the issue.
Lastly, the current setup, which captures images at zero yaw with a 90-degree tilt, limits the range of available angles. However, this constraint presents an opportunity for future enhancement. Future versions could incorporate more advanced algorithms and utilize APIs beyond the static Google Maps API to dynamically adjust both tilt and yaw. This would expand the range of angles available, offering greater flexibility and more comprehensive data capture capabilities.
@@ 5 Conclusion
SkyAI Sim’s functionality allows users to customize various parameters, including flight altitude, camera aspect ratio, diagonal field of view, and image overlap. This flexibility supports a wide range of applications, from capturing low-altitude images for simple tasks to generating comprehensive datasets for complex projects such as deep learning. In addition to its versatility, SkyAI Sim’s open-source nature encourages further development and adaptation to other missions beyond its initial design. The tool’s ability to produce extensive datasets without the need for physical UAV deployment represents a significant advancement in dataset generation for VBN algorithm developments. SkyAI Sim presents a compelling solution to the challenges of aerial image capture, offering a powerful and flexible tool for a variety of applications, and opening new avenues for research and development in VBN and beyond.

### v2

@@ 3.3.1 Input Formats for the Argument Parser
    The default is "TerrAInav". It can be switched to "VBN" to access and process the Supplementary data.
@@ 3.4 Post Processing
To cleanup the data as an optional step, a threshold entropy is defined. The threshold is selected according to the distribution of entropies in all images, along with experimental observations provided in Figure 6. Specifically in the Agricenter dataset, the images below an entropy of 2.1 lack significant details. To confirm this value’s reliability and ensure it doesn’t filter out substantial valuable data, a barplot and histogram of entropies for all available roadmap images are shown in Figures 9 and 10. This ensures most data is above the threshold, making it safe to discard images below $2.1$. TerrAInav performs data cleaning based on the map type. If the map type is "satellite" or anything other than "roadmap", it first checks for "roadmap" data in the data directory. If available, images are filtered according to "roadmap" images with an entropy above the default threshold (2.1). If no roadmap data exists, images are filtered based on their original map type.
@@ 4 Discussion
TerrAInav Sim opens the door to numerous applications, but it does have limitations that could affect its utility in certain scenarios. One notable limitation arises from the reliance on the Google Maps Static API, which prevents the determination of the exact date of the imagery. This can lead to discrepancies in temporal accuracy, especially in applications where precise timestamping is crucial, such as change detection in environmental monitoring or tracking land use over time in agricultural applications. This limitation, while not critical for most of TerrAInav Sim’s potential applications, is still an important area for future improvement. To address this limitation, a potential solution is to integrate additional data sources that provide time-stamped satellite imagery. For instance, services like Google Earth Engine [3] offer access to historical satellite imagery with precise timestamps, which could be leveraged through their APIs. However, their archive still has their own limitations. Integrating these services into TerrAInav Sim would significantly improve the temporal accuracy of the imagery, making the tool more suitable for applications that require historical comparisons, event tracking, or temporal analysis.
Another limitation lies in the idealized nature of the simulation, which does not account for the complexities and inaccuracies typical of real-world drone missions. Factors such as sensor faults, varying environmental conditions, and obstacles like buildings or trees can significantly impact the quality of images captured during actual operations. For example, in tasks like urban mapping or vegetation analysis, these inaccuracies could introduce errors in feature extraction or object recognition. Currently, TerrAInav Sim provides a controlled, distortion-free environment, but this does not reflect the unpredictability of real-world operations. To mitigate this, incorporating error models simulating sensor noise, dynamic obstacles, and environmental factors (e.g., wind or lighting changes) could better represent real-world conditions. This would allow users to assess how their systems would perform in diverse, unpredictable scenarios, which is particularly important for applications like environmental monitoring or disaster response.
The current simulator also uses a high-altitude image capture and cropping method to simulate low-altitude drone behavior. While the proposed approach provides a reasonable approximation, it does not replicate the distortions and misalignments experienced during actual low-altitude flights. This discrepancy can be particularly problematic in applications requiring precise 3D modeling or when operating in densely built-up areas where buildings can create severe perspective distortions. To address this, future versions of the simulator could implement custom distortion models that account for terrain elevation, building height, and the specific angle of capture. This would allow for more realistic simulations of low-altitude operations, improving the accuracy of applications like urban modeling or emergency response planning.
Furthermore, there is a trade-off between image quality and processing speed. High-quality images require more computational resources, which can increase processing time, potentially limiting the simulator’s utility in real-time applications, such as autonomous navigation or obstacle avoidance in dynamic environments. On the other hand, faster image acquisition and processing can lead to a reduction in image fidelity, which may compromise the accuracy of object detection and map generation. In future versions, a dynamic adjustment system that optimizes the balance between image quality and processing speed, perhaps based on the specific needs of the task at hand (e.g., prioritizing speed for real-time navigation vs. quality for detailed mapping), could provide a more flexible and practical solution for users.
Lastly, the current image capture setup, which limits yaw to zero and tilt to 90 degrees, restricts the available angles for data capture. This limitation is particularly restrictive in applications that require a full 360-degree view for accurate environmental modeling or surveillance. Expanding the simulator to include dynamic adjustments for both yaw and tilt would allow for a broader range of angles, enabling more comprehensive data capture for tasks like infrastructure inspection, where capturing multiple angles of a structure is essential for thorough analysis.
@@ 5 Conclusion
TerrAInav Sim’s functionality allows users to customize various parameters, including flight altitude, camera aspect ratio, diagonal field of view, and image overlap. This flexibility supports a wide range of applications, from capturing low-altitude images for simple tasks to generating comprehensive datasets for complex projects. The ability to quickly and efficiently capture large volumes of images while utilizing minimal memory makes it an ideal candidate for deep learning. Its potential applications span various fields, including classification, pattern recognition, localization, and image-to-image translation, showcasing just a few of its capabilities in deep learning. By offering a faster and simpler alternative to commercial flight simulators, TerrAInav Sim accelerates experimentation and validation processes, making it an invaluable resource for future advancements in UAV-based geospatial analysis, environmental monitoring, and autonomous navigation research.
In addition to its versatility, TerrAInav Sim’s open-source nature encourages further development and adaptation to other missions beyond its initial design. The tool’s ability to produce extensive datasets without the need for physical UAV deployment represents a significant advancement in dataset generation for VBN algorithm developments. TerrAInav Sim presents a compelling solution to the challenges of aerial image capture, offering a powerful and flexible tool for a variety of applications, and opening new avenues for research and development in VBN and beyond.
By providing a customizable, cost-effective tool for generating diverse image datasets, TerrAInav Sim lowers barriers to entry for research in autonomous navigation, environmental monitoring, and urban planning. Its open-source nature encourages continued development, making it a valuable asset for advancing machine learning and remote sensing applications.